OpenStudy (anonymous):

Find the quotient of z1 by z2. Express your answer in trigonometric form. z1 = 4(cos(pi/3) + isin(pi/3)) z2 = 3(cos(2pi/5) + isin(2pi/5)) A. 4/3(cos(-pi/15) + isin(-pi/15)) B. 12(cos(-pi/15) + isin(-11pi/15)) C. 12(cos(11pi/15) + isin(11pi/15)) D. 4/3(cos(11pi/15) + isin(11pi/15))

OpenStudy (anonymous):

@hartnn

Parth (parthkohli):

What tells me that you have the Euler's Formula involved?

OpenStudy (anonymous):

the what? :/

Parth (parthkohli):

Never mind. Actually no, it's pretty simple:\[\large e^{ix}=\cos(x) + i\sin(x)\]

OpenStudy (anonymous):

how does that help me?

OpenStudy (anonymous):

ello..?

Parth (parthkohli):

Not sure enough.

OpenStudy (anonymous):

say what

OpenStudy (anonymous):

@ZeHanz perhaps you can help?

OpenStudy (foolaroundmath):

\[\Large z_{1} = 4e^{i \pi/3}, z_{2} = 3e^{i2\pi/5}\] \[\Large z_{1}/z_{2} = \frac{4e^{i \pi/3}}{3e^{i2\pi/5}} = \frac{4}{3}e^{i(\pi/3 - 2\pi/5)}\] \[\Large = \frac{4}{3}e^{-i\pi/15} = \frac{4}{3}(\cos{\frac{-\pi}{15}}+i\sin{\frac{-\pi}{15}})\] Hence (A)

OpenStudy (zehanz):

@Spectrum: I think @FoolAroundMath did a pretty good job!

OpenStudy (anonymous):

well i tbh would have considered it foolish to accept that answer with out letting you have your say and it said you were still typing ..but ty

OpenStudy (zehanz):

Thank you for that!